OpenStudy (anonymous):

Please solve -24c^2=36c

OpenStudy (anonymous):

-24c^2 - 36c = 0 12c(-2c - 3) = 0 12c = 0 c = 0 -2c - 3 = 0 -2c = 3 c = -3/2 c = 0, -3/2